# AGN-heated dust revealed in “Little Red Dots”

### v1

@@ B.2 Field-to-field variations
An important aspect that our analysis relies on is the combination of multiple NIRCam and MIRI survey fields. While this is necessary to improve the statistics and the stacked S/N, it might potentially introduce systematics related to different JWST coverage and depth, which could affect the properties of photometrically-selected LRDs. We thus verify that our stacking results are robust against field-to-field variations, by repeating the same analysis separately for each field. The second panel in Fig. 5 displays the median stacked SED obtained in each field (represented by different filled symbols) overlaid to our full stacking results (grey open circles). For visual purposes, here we relax the detection threshold to S/N2 (hence upper limit fluxes are set to twice the noise level if S/N2). Despite their heterogeneous spectral coverage across MIRI bands, all single-field median SEDs are consistent with the combined median SED. Moreover, we verify that stacking the same exact 22 objects having NIRCam/MIRI coverage in all bands (from the GOODS-S field), the resulting SED is fully consistent with that obtained via stacking the full LRD sample. These checks strengthen that similarly-selected LRDs in different fields and with heterogeneous spectral coverage share a common median SED.
B.3 Effect of photometric redshifts
@@
A relevant caveat in our analysis is the photometric selection of LRDs. Specifically, photometric redshift uncertainties can be notably high, potentially biasing the LRD selection. We iterate that for the spectroscopic subsample (85/302 objects) there is good agreement between photometric and spectroscopic redshifts (Sect. 2). However, we test this directly by stacking NIRCam and MIRI images solely for the spectroscopic subsample. The third panel in Fig. 5 displays the corresponding median stacked SED (blue filled circles). Numbers are given in Table 1 for each field: when combining all fields there are between 21 and 47 objects in MIRI (except for f560w, f1280w and f2550w with 5 objects due to a smaller areal coverage) and 85 objects in NIRCam. Despite the S/N being smaller than that of the full sample, the two median stacked SEDs agree remarkably well with each other.
B.4 Effect of LRD selection
In the literature, LRDs are inevitably selected with distinct criteria, depending on the availability of multi-band NIRCam and MIRI imaging, and rest-frame optical spectra. While our LRD sample meets the same compactness and V-shape criteria across all fields, we compare our stacked SED against other LRD SEDs based on different selections. The fourth panel of Fig. 5 shows for comparison: the median stacked SED from Akins et al. (2024; green symbols); the corresponding galaxy and AGN best-fit LRD SED from Casey et al. (2024; dot-dashed lines); the ALMA upper limit from Casey et al. (2025; downward triangle) and the observed CAPER-z9 NIRSpec spectrum (at =9.288; Taylor et al. 2025) shifted to =6.2 (orange symbols).
Akins et al. (2024) selected 434 LRDs in the COSMOS-Web area (0.5 deg2) that simultaneously meet a compactness (f444w) cut, a very red colour cut (f444w/f277w) and a bad brown drarf fit, without explicitly requiring a “V-shaped” SED due to limitation in NIRCam imaging. Their median SED agrees remarkably well with ours, though their statistics in MIRI is limited to 7.7m (over 1/3 of COSMOS-Web) and 18m (from the inner PRIMER-COSMOS footprint, 4% of the full sample), this latter leading to tentative detection (0.430.15 mJy). We show that their 18m flux is fully consistent with the rising MIR slope seen longward 10m. Their upper limits are here re-scaled to 3 (Herschel and SCUBA limits are too shallow to enter the y-scale). Despite the different LRD selection and survey field (25% of our LRDs are in Akins et al. 2024), this agreement further suggests that the compactness and the conservative colour selection by Akins et al. (2024) broadly compensate the need for a V-shaped SED, leading to self-consistent LRD samples.
In the same fourth panel of Fig. 5, we compare our stacks against the median LRD SED by Casey et al. (2024), that is inferred from a joint LRD sample (675 LRDs) from Akins et al. (2024) and Kokorev et al. (2024). This latter consists of a compactness (in f444w) criterion and various colour cuts that mimic a V-shape requirement (see their Sect.3.1 in Kokorev et al. 2024 for details). In their work, Casey et al. (2024) performed stacking in rest-frame wavelength bins, following the observed SED of each LRD. The median stacked SED was then redshifted to =6.2 (as for our sample) and fitted with either a pure AGN (red) or a pure galaxy template (blue). We emphasize that our median stacks are fully consistent with their median AGN SED within their uncertainties (here not reported). If any, the AGN SED from Casey et al. (2024) stands slightly higher than our MIRI points, that partly reflects the MIRI upper limits in their study (20 objects per bin at 0.8m). Our stacks and their SED are also mutually consistent with the ALMA upper limit derived in Casey et al. (2025) by stacking 60 LRDs (among the Akins et al. 2024 sample) at =1.3 mm.
In addition, we compare our stacks against the NIRSpec spectrum of CAPER-z9 at =9.288 (Taylor et al. 2025), which is complemented by MIRI data at 10m (detection) and 21m (upper limit). This continuous spectrum enables a direct comparison with our SED shape over a broad wavelength range. After moving the CAPER-z9 spectrum and photometry to =6.2 (orange symbols), we again find a reassuring consistency with our data.
Another meaningful comparison with MIRI studies of LRDs is against the sample of Williams et al. (2024), who stacked LRDs in all MIRI bands within the GOODS-SMILES footprint. They found a flat IR-SED up to 18m (and upper limits longward), thus excluding a dominant contribution from AGN-heated dust. Besides the small numbers (only 9 objects) and the slightly higher redshift compared to our sample (7.7 instead of 6.2), their LRDs were pre-selected from a pool of H-band dropouts (i.e. HST-dark galaxies), hence likely biased towards the most obscured sources. We repeat the stacking for the same 9 sources, finding the same median SED as in Williams et al. (2024). This check suggests that their sample might be sensitive to more obscured sources than conventional LRDs, especially considering their H-dropout selection, combined with the lack of compactness and V-shaped requirements. These differences might link to a different obscuration origin relative to typical LRDs. Hence we caution that these 9 LRDs are not representative of our photometrically-selected LRDs (from Kocevski et al. 2025), albeit they are certainly an interesting sub-sample to follow-up.
B.5 Effect of a broad redshift range
@@
As shown in Fig. 4, our targets span a very broad redshift range, from 2 to 11, with over 90% of the sample being within the range (48). Because we perform stacking in the observed frame, the contribution to each stacked point comes from very different parts of the rest-frame spectrum. As mentioned in Sect. 4.1, we already account for this by convolving the filter response with the underlying redshift distribution. However, this approach assumes that each galaxy contributes evenly to the stack, regardless of its redshift, whereas the flux dimming can vary significantly across the full range. We thus test the impact of the redshift range to the stacked fluxes by splitting our LRDs into two similarly-populated -bins at 46 (126 LRDs) and 68 (121 LRDs). We show this comparison in the bottom panel of Fig. 5. The lower- bin shows slightly brighter fluxes and steeper MIRI slope compared to the higher- bin, as expected given the smaller distance and the longer rest-frame wavelength probed (e.g. 3.5 m instead of 2.5 m at MIRI-21 m), respectively. Overall, this check suggests that LRDs selected at different redshifts have broadly-consistent stacked SEDs, with no evidence for a galaxy-like drop at rest 1–2 m. We further note that emission lines are unlikely to systematically boost the median stacks, as this would be the case if a large fraction of sources were clustered in a narrow redshift range (such that a rest-frame emission line would systematically fall into a given filter). Given the broad similarity of the two SEDs in two different -bins, we believe this is unlikely the case.
Appendix C Input SED-fitting parameters
@@
[…]
@@
Appendix D Output SED-fitting parameters

### v2

@@ A.3 Field-to-field variations
An important aspect that our analysis relies on is the combination of multiple NIRCam and MIRI survey fields. While this is necessary to improve the statistics and the stacked S/N, it might potentially introduce systematics related to different JWST coverage and depth, which could affect the properties of photometrically-selected LRDs. We thus verify that our stacking results are robust against field-to-field variations, by repeating the same analysis separately for each field. The third panel in Fig. 6 displays the median stacked SED obtained in each field (represented by different filled symbols) overlaid to our full stacking results (grey open circles). For visual purposes, here we relax the detection threshold to S/N2 (hence upper limit fluxes are set to twice the noise level if S/N2). Despite their heterogeneous spectral coverage across MIRI bands, all single-field median SEDs are consistent with the combined median SED. Moreover, we verify that stacking the same exact 22 objects having simultaneously NIRCam and MIRI coverage in all bands (from the GOODS-S field), the resulting SED is fully consistent with that obtained via stacking the full LRD sample. These checks strengthen that similarly-selected LRDs across different fields share a common median SED, despite their and heterogeneous NIRCam/MIRI spectral coverage.
A.4 Effect of photometric redshifts
@@
A relevant caveat in our analysis is the photometric selection of LRDs. Specifically, photometric redshift uncertainties can be notably high, potentially biasing the LRD selection. We iterate that for the spectroscopic subsample (96/302 objects) there is good agreement between photometric and spectroscopic redshifts (Sect. 2). However, we test this directly by stacking NIRCam and MIRI images solely for the spectroscopic subsample. The fourth panel in Fig. 6 displays the corresponding median stacked SED (blue filled circles). Numbers are given in Table 1 for each field: when combining all fields there are between 36 and 53 objects in MIRI (except for f560w, f1280w and f2550w with 5 objects due to a smaller areal coverage) and 96 objects in NIRCam. Despite the S/N being smaller than that of the full sample, the two median stacked SEDs agree remarkably well with each other.
We also explore whether a cleaner selection of “bona fide” AGN based on broad emission lines yields similar trends. From the NIRSpec Merged Table of the DJA, we isolate 57/96 (59%) objects with broad Hα or Hβ emission lines, as inferred from the best-fit continuum fluxes, line intensity and equivalent widths. This fraction agrees with that reported in the literature (e.g. Greene et al. 2024; Hviding et al. 2025). The corresponding median stacked SED of BL AGN (empty starred symbols), displays no significant difference with respect to the full spec- (96) subsample, albeit a lower S/N in all bands due to smaller statistics.
A.5 Effect of a broad redshift range
@@
As shown in Fig. 1, our targets span a very broad redshift range, from 2 to 11, with over 90% of the sample being within the range (48). Because we perform stacking in the observed frame, the contribution to each stacked point comes from very different parts of the rest-frame spectrum. As mentioned in Sect. 4.1, we already account for this by convolving the filter response with the underlying redshift distribution. However, this approach assumes that each galaxy contributes evenly to the stack, regardless of its redshift, whereas the flux dimming can vary significantly across the full range. We thus test the impact of the redshift range to the stacked fluxes by splitting our LRDs into two similarly-populated -bins at 46 (126 LRDs) and 68 (121 LRDs). We show this comparison in the bottom panel of Fig. 6. The lower- bin shows slightly brighter fluxes and steeper MIRI slope compared to the higher- bin, as expected given the smaller distance and the longer rest-frame wavelength probed (e.g. 3.5 m instead of 2.5 m at MIRI-21 m), respectively. We also note that fluxes are in energy units, hence the apparent flattening of the 68 NIR SED translates into a monothonic rise in flux density () units, as seen in the full sample. Overall, this check suggests that LRDs selected at different redshifts have broadly-consistent stacked SEDs, with no evidence for a galaxy-like drop at rest 1–2 m. We further note that emission lines are unlikely to systematically boost the median stacks, as this would be the case if a large fraction of sources were clustered in a narrow redshift range (such that a rest-frame emission line would systematically fall into a given filter). Given the broad similarity of the two SEDs in two different -bins, we believe this is unlikely the case.
Appendix B Input SED-fitting parameters
@@
| Single stellar population | Initial mass function | – | Chabrier (2003) |
@@
Appendix C Output SED-fitting parameters